Solve the equation and state any extraneous solutions. Thanks so much!
5x/ x + 1 + 1/ x = 5

Answer by stanbon(75887)

5x/ (x + 1) + 1/ x = 5
Multiply thru by x(x+1) to get:
5x^2+(x+1)=5x(x+1)
5x^2+x+1=5x^2+5x
4x=1
x=1/4